# Schwarzschild black hole surrounded by a cloud of strings in the background of perfect fluid dark matter

### v1

@@ 1 Introduction
The manuscript is organized as follows: Section 2 provides an overview of the mathematical framework for the SBH in the presence of a CoS and PFDM. Section 3 discusses the black hole’s thermodynamic properties, highlighting the effects of the string and PFDM parameters. Section 4 explores the geodesic structure, followed by Section 5, which examines the black hole’s shadow in the context of a CoS and PFDM. Section 6 investigates the quasinormal modes of the system. Finally, the manuscript concludes with a summary of the key findings.
@@ 2 SBH immersed in a CoS and PFDM
To examine the impact of the CoS and PFDM on the black hole, we plot the metric function versus the radius diagrams in Fig.1.
@@
In both cases, we observe that increasing the CoS parameter results in a downward shift of the metric function, with the effect being more pronounced near the black hole horizon. In the case of $\alpha=0.8$, see Fig. 1, the metric function crosses into positive values more quickly, implying a stabilizing effect of the CoS parameter and pushing the horizon outward. Conversely, for $\alpha=-0.8$, see Fig. 1, the metric function starts at more negative values and remains below zero for larger $r$, indicating that the presence of a negative PFDM parameter further reduces the spacetime curvature and shifts the horizon inward. As $r$ increases, the metric function approaches a constant value in both cases, with the influence of $a$ and $\alpha$ diminishing at larger radii.
@@ 3 Thermodynamics
$\kappa=-\left.\frac{1}{2\sqrt{-g_{00}g_{11}}}\frac{\partial g_{00}}{\partial r}\right|_{r=r_{H}}.$ (18)
@@
Fig. 3 illustrates the behavior of heat capacity with the positive PFDM parameter. The results show that the heat capacity is consistently negative, which signifies the thermodynamic instability of the black hole. Notably, the curves for varying values of $a$ are distinct, particularly for larger values of $r_{H}$.
In contrast, Fig. 3 depicts a negative PFDM parameter. Here, we observe that the heat capacity reaches zero at $r_{H}=r_{min}$, which marks the minimum size at which the black hole ceases to radiate. The stable black hole is characterized by a positive heat capacity within the range $r_{min}<r_{H}<r_{crt}$, indicative of a smaller black hole. Conversely, for $r_{H}>r_{crt}$ the system becomes unstable, exhibiting a negative heat capacity associated with a larger, more massive black hole. Additionally, the stable range $r_{min}<r_{H}<r_{crt}$ where the heat capacity remains positive expands as $a$ increases. This observation suggests that smaller values of $a$ result in a quicker cessation of radiation and evaporation for the black hole.
@@
In Fig. 4, we illustrate the variation of the free energy with respect to $r_{H}$. We observe that the Gibbs free energy reaches a minimum at
$r_{H}=-\frac{2\alpha}{1-a},$ (31)
@@ 4.1 Timelike geodesics
$\frac{d^{2}}{dr^{2}}V_{\mathrm{eff}}\left(r\right)=0,$ (48)
@@
$02rf"\left(r\right)f\left(r\right)+6f\left(r\right)f^{\prime}\left(r\right)-4rf^{\prime}\left(r\right)^{2}=$ 0 . (49)
@@ 4.2 Null geodesics
$\left.\frac{d}{dr}V_{\mathrm{e}}\right|_{r=r_{\rm{ph}}}$ $=$ $0$ (52)
@@
M87 $\left(6.5\pm 0.90\right)\times 10^{9}$ $42\pm 3$ $16800$
Table 3: Observational measurements of Sgr. A and M87 black holes.
@@
We follow the bounds provided in [106]. For Sgr A the permitted range is
$4.55\leq\frac{R_{\rm{S}}}{M}\leq 5.22,$ (61)
while for M87, the allowed range is
$4.31\leq\frac{R_{S}}{M}\leq 6.08$ (62)
Our results are presented in Fig. 10, which includes the confidence intervals.
(a) Sgr. A.
(b) M87.
@@
Although the constraints on $a$ vary for Sgr A and M87, the bounds remain relatively similar upon comparison.
@@ 5.1 WKB method
$\psi\left(t,r,\theta,\varphi\right)=e^{-i\omega t}\frac{\Psi_{\omega,L}\left(r\right)}{r}Y_{L,\mu}\left(\theta,\varphi\right),$ (64)
where $\omega$ represents the frequency, and $Y_{L,\mu}\left(\theta,\varphi\right)$ are the spherical harmonics. Substituting this decomposition into Eq. (63) transforms it into a Schrödinger-like equation:
@@
$\mathcal{V}\left(r\right)=\frac{f\left(r\right)}{r}\frac{df\left(r\right)}{dr}+\frac{f\left(r\right)L\left(L-1\right)}{r^{2}}.$ (66)
@@
where $V_{0}$ corresponds to the height of the effective potential, and $V_{0}"$ denotes the second derivative of the potential with respect to the tortoise coordinate, evaluated at its maximum $r_{0}^{\ast}$, respectively. The term $\Lambda_{i}$ accounts for corrections from higher-order terms in the WKB approximation, while $n=0,1,2,...$, is the overtone number.
Now, we calculate the QNM frequencies of the SBH by applying the 6th-order WKB approximation method for different angular quantum numbers $L$ and overtone numbers $n$ in the context of a massless scalar field. These frequencies are separated into real and imaginary components, where the real part refers to the oscillation frequency and the imaginary part corresponds to the decay rate. We tabulate the QNMs in Table 5.
@@
$L$ $n$ $\omega_{\rm{WKB}}(a=0.2)$ $\omega_{\rm{WKB}}(a=0.5)$ $\omega_{\rm{WKB}}(a=0.2)$ $\omega_{\rm{WKB}}(a=0.5)$
@@
We observe that the imaginary components of all frequencies are negative, confirming the stability of the black hole. Furthermore, these imaginary components describe the decay rate of the oscillations, which is tied to the dynamical evolution of the QNMs.
@@
Figure 11: The quasinormal modes of the black holes in scalar field for the state $L=7$, $\alpha=0.5$ and $M=1$
It is clear that for a fixed $n$, both $\omega_{\rm{R}}$ and $\omega_{\rm{I}}$ decrease as $a$ increases. This suggests that a reduction in $a$ results in an increase in the oscillation frequency and a faster decay rate. Next, by setting $a=0.4$ we examine the impact of PFDM on the QNMs, as depicted in Fig. 12.
Figure 12: QNMs of the SBH in the scalar field for the state $L=7$, $a=0.4$ and $M=1$.
@@ 5.2 Connection between shadow and quasinormal modes
$\omega_{\rm{S}}=\omega_{\rm{R}}-i\omega_{\rm{I}}=\Omega L-i\lambda\left(n+\frac{1}{2}\right),$ (69)
@@
$\lambda=\left.\sqrt{-\frac{V"\left(r\right)}{2\dot{t}^{2}}}\right|_{r=r_{\rm{ph}}}=\sqrt{\frac{\left(2f\left(r_{\rm{ph}}\right)-r_{\rm{ph}}^{2}f^{\prime\prime}\left(r_{\rm{ph}}\right)\right)f\left(r_{\rm{ph}}\right)}{2r_{\rm{ph}}^{2}}},$ (71)
@@
$L$ $\omega_{\rm{WKB}}$ $\omega_{\rm{S}}$ $\Delta_{\rm{R}}\%$ $\Delta_{\rm{I}}\%$
@@
$L$ $\omega_{\rm{WKB}}$ $\omega_{\rm{S}}$ $\Delta_{\rm{R}}\%$ $\Delta_{\rm{I}}\%$
@@
$L$ $\omega_{\rm{WKB}}$ $\omega_{\rm{S}}$ $\Delta_{\rm{R}}\%$ $\Delta_{\rm{I}}\%$
@@
$L$ $\omega_{\rm{WKB}}$ $\omega_{\rm{S}}$ $\Delta_{\rm{R}}\%$ $\Delta_{\rm{I}}\%$
@@
We observe that as $L$ increases, the differences in both the real and imaginary components of the frequency gradually diminish. This trend suggests that the eikonal limit method is highly accurate for large angular momentum values, producing results that closely align with the WKB approximation. This finding holds significant practical value, as the eikonal limit method is generally more straightforward for analytical computations. Furthermore, high-precision results provide reliable data for investigating the stability and oscillatory behavior of black holes and other compact objects.
@@ Acknowledgments
B. C. L. is grateful to Excellence project PřF UHK 2211/2023-2024 for the financial support.

### v2

@@ Abstract
Keywords: Clouds of strings; perfect fluid dark matter; Schwarzschild black hole; thermodynamics; geodesics; shadows; quasinormal modes.
@@ 1 Introduction
The manuscript is organized as follows: Section 2 provides an overview of the mathematical framework for the SBH in the presence of a CoS and PFDM. Section 3 discusses the black hole’s thermodynamic properties, highlighting the effects of the string and PFDM parameters. Section 4 explores the geodesic structure, followed by Section 5, which examines the black hole’s shadow in the context of a CoS and PFDM. Section 6 investigates the quasinormal modes of the system. Finally, the manuscript concludes with a summary of the key findings. It is important to note that throughout the manuscript, natural units are adopted, where $\hbar=k_{B}=G=c=1$.
@@ 2 SBH immersed in a CoS and PFDM
To examine the impact of the CoS and PFDM on the black hole, we plot the metric function versus the radius diagrams in Fig.1. Here, we consider both positive and negative values of $\alpha$ in the remainder of our analysis to explore the full range of possible effects on black hole properties.
@@
In both scenarios, increasing the CoS parameter causes a downward shift in the metric function, thereby influencing the horizon’s location. For $\alpha=0.8$, as the CoS parameter increases, the metric function reaches positive values at progressively larger radial distances, effectively moving the horizon outward. In contrast, a distinct structural configuration emerges in the scenario with $\alpha=-0.8$. As shown in Fig. 1, the metric function displays two distinct horizons: event horizon $r_{H}$ and an inner Cauchy horizon $r_{C}$, in contrast to the single horizon observed in Fig. 1. This result indicates that the PFDM parameter has a substantial impact on spacetime curvature, and consequently, on the black hole’s thermal properties, particularly in relation to its stability. Finally, as $r$ increases, the metric function approaches a constant value in both cases, with the influence of $a$ and $\alpha$ diminishing at larger radii.
Upon comparison with the existing literature, we observe similar results. For instance, Al-Badawi et al. [72], who investigated the GUP-corrected black holes in the presence of a CoS, found that increasing the CoS parameter values leads to a larger horizon radius, a trend that also holds in our study for both positive and negative PFDM scenarios.
@@ 3 Thermodynamics
$\kappa=-\left.\frac{1}{2\sqrt{-g_{00}g_{11}}}g_{00}^{{\prime}}\right|_{r=r_{H}}.$ (18)
@@
Fig. 3 illustrates the behavior of heat capacity with the positive PFDM parameter. The results show that the heat capacity is consistently negative, which signifies the thermodynamic instability of the black hole. Notably, the curves for varying values of $a$ are distinct, particularly for larger values of $r_{H}$. In contrast, Fig. 3 depicts the heat capacity with a negative PFDM parameter, where we observe that the heat capacity reaches zero at $r_{H}=r_{min}$, marking the minimum size at which the black hole ceases to radiate. The stable black hole is characterized by a positive heat capacity within the range $r_{min}<r_{H}<r_{crt}$, indicative of a smaller black hole. Conversely, for $r_{H}>r_{crt}$ the system becomes unstable, exhibiting a negative heat capacity associated with a larger, more massive black hole. Additionally, the stable range $r_{min}<r_{H}<r_{crt}$ where the heat capacity remains positive expands as $a$ increases. This observation suggests that smaller values of $a$ result in a quicker cessation of radiation and evaporation for the black hole.
In comparison to existing literature, such as Chaudhary et al. [73], which demonstrates that the CoS induces critical regions where black holes transition between stability and instability under EGUP corrections, our findings further highlight that the CoS, in conjunction with PFDM, significantly modifies the thermodynamic properties and stability thresholds of black holes.
@@
In Fig. 4, we illustrate the variation of the free energy with respect to $r_{H}$.
@@
We observe that the Gibbs free energy reaches a minimum at
$r_{H}=-\frac{2\alpha}{1-a}{=r_{crt},}$ (31)
@@ 4.1 Timelike geodesics
$\frac{d^{2}}{dr^{2}}V_{\mathrm{eff}}\left(r\right)=f^{\prime\prime}\left(r\right)\left(1+\frac{L^{2}}{r^{2}}\right)-f^{\prime}\left(r\right)\frac{4L^{2}}{r^{3}}+f\left(r\right)\frac{6L^{2}}{r^{4}}=0.$ (48)
@@
$02rf^{\prime\prime}\left(r\right)f\left(r\right)+6f\left(r\right)f^{\prime}\left(r\right)-4rf^{\prime}\left(r\right)^{2}=$ 0 . (49)
@@ 4.2 Null geodesics
$\left.V_{\mathrm{e}}^{\prime}\right|_{r=r_{\rm{ph}}}$ $=$ $0$ (52)
@@
Table 3: Observational measurements of Sgr. A black hole.
@@
It is worth noting that our results indicate an inverse proportionality between the photon sphere radius and $(1-a)$, a behavior also observed in the literature. In particular, in [72] Al-Badawi et al. demonstrated that both the photon sphere radius and the impact factor, influenced by the presence of the CoS, exhibit the same inverse scaling with $(1-a)$.
Then, we follow the bounds provided in [106]. For Sgr A the permitted range is
$4.55\leq\frac{R_{\rm{S}}}{M}\leq 5.22$ (61)
Our result is presented in Fig. 10, which includes the confidence interval.
@@ 5.1 WKB method
$\psi\left(t,r,\theta,\varphi\right)=e^{-i\omega t}\frac{\Psi_{\omega,{\ell}}\left(r\right)}{r}Y_{{\ell},\mu}\left(\theta,\varphi\right),$ (63)
where $\omega$ represents the frequency, $\ell$ is the angular quantum number, and $Y_{{\ell},\mu}\left(\theta,\varphi\right)$ are the spherical harmonics. Substituting this decomposition into Eq. (62) transforms it into a Schrödinger-like equation:
@@
$\mathcal{V}\left(r\right)=\frac{f\left(r\right)}{r}f^{\prime}(r)+\frac{f\left(r\right){\ell\left(\ell-1\right)}}{r^{2}}.$ (65)
@@
where $V_{0}$ corresponds to the height of the effective potential, and $V_{0}^{\prime\prime}$ denotes the second derivative of the potential with respect to the tortoise coordinate, evaluated at its maximum $r_{0}^{\ast}$, respectively. The term $\Lambda_{i}$ accounts for corrections from higher-order terms in the WKB approximation, while $n=0,1,2,...$, is the overtone number.
Now, we calculate the QNM frequencies of the SBH by applying the 6th-order WKB approximation method for different angular quantum numbers $\ell$ and overtone numbers $n$ in the context of a massless scalar field. These frequencies are separated into real and imaginary components, where the real part refers to the oscillation frequency and the imaginary part corresponds to the decay rate. We tabulate the QNMs in Table 5.
@@
${\ell}$ $n$ $\omega_{\rm{WKB}}(a=0.2)$ $\omega_{\rm{WKB}}(a=0.5)$ $\omega_{\rm{WKB}}(a=0.2)$ $\omega_{\rm{WKB}}(a=0.5)$
@@
We observe that the imaginary components of all frequencies are negative, confirming the stability of the black hole. This also indicates the damping behavior of the oscillations, which is tied to the dynamical evolution of the QNMs.
@@
Figure 11: The quasinormal modes of the black holes in scalar field for the state $\ell=7$, $\alpha=0.5$ and $M=1$.
It is clear that for a fixed $n$, both $\omega_{\rm{R}}$ and $\omega_{\rm{I}}$ decrease as $a$ increases. This suggests that a reduction in $a$ results in an increase in the oscillation frequency and a faster decay rate. This result is consistent with the findings presented by Al-Badawi et al. in Table 2 of [72].
Next, by setting $a=0.4$ we examine the impact of PFDM on the QNMs, as depicted in Fig. 12.
Figure 12: QNMs of the SBH in the scalar field for the state $\ell=7$, $a=0.4$ and $M=1$.
@@ 5.2 Connection between shadow and quasinormal modes
$\omega_{\rm{S}}=\omega_{\rm{R}}-i\omega_{\rm{I}}=\Omega{\ell}-i\lambda\left(n+\frac{1}{2}\right),$ (68)
@@
$\lambda=\left.\sqrt{-\frac{V^{\prime\prime}\left(r\right)}{2\dot{t}^{2}}}\right|_{r=r_{\rm{ph}}}=\sqrt{\frac{\left(2f\left(r_{\rm{ph}}\right)-r_{\rm{ph}}^{2}f^{\prime\prime}\left(r_{\rm{ph}}\right)\right)f\left(r_{\rm{ph}}\right)}{2r_{\rm{ph}}^{2}}},$ (70)
@@
$\ell$ $\omega_{\rm{WKB}}$ $\omega_{\rm{S}}$ $\Delta_{\rm{R}}\%$ $\Delta_{\rm{I}}\%$
@@
$\ell$ $\omega_{\rm{WKB}}$ $\omega_{\rm{S}}$ $\Delta_{\rm{R}}\%$ $\Delta_{\rm{I}}\%$
@@
$\ell$ $\omega_{\rm{WKB}}$ $\omega_{\rm{S}}$ $\Delta_{\rm{R}}\%$ $\Delta_{\rm{I}}\%$
@@
$\ell$ $\omega_{\rm{WKB}}$ $\omega_{\rm{S}}$ $\Delta_{\rm{R}}\%$ $\Delta_{\rm{I}}\%$
@@
We observe that as $\ell$ increases, the differences in both the real and imaginary components of the frequency gradually diminish. This trend suggests that the eikonal limit method is highly accurate for large angular momentum values, producing results that closely align with the WKB approximation. This finding holds significant practical value, as the eikonal limit method is generally more straightforward for analytical computations. Furthermore, high-precision results provide reliable data for investigating the stability and oscillatory behavior of black holes and other compact objects.
@@ 6 Conclusion
Finally, it is important to emphasize that black holes should not be regarded as isolated entities. In nature, they constantly interact with their surroundings and are rarely unaffected by neighboring fields. Consequently, they are typically in a perturbed state due to the influence of other fields in their vicinity. The presence of PFDM and CoS around black holes offers valuable opportunities to test and expand gravity theories. PFDM modifies black hole thermodynamics and particle orbits, influencing key properties such as temperature, phase transitions, and gravitational lensing. These effects establish a bridge between the cosmological role of dark matter and its impact on black hole physics. On the other hand, CoS affects the geometry of spacetime and the physical phenomena nearby. The mathematical structure of string theory has consistently demonstrated that it encompasses all the essential components for a quantum theory of gravity, including valuable insights into black hole entropy. However, string theory is not merely a theory of quantum gravity; it also incorporates particles and interactions, providing the tools necessary to address contemporary questions in cosmology. PFDM specifically addresses the coupling between dark matter and gravity, while the effects of CoS provide concrete tests of string theory-inspired modifications and potential signatures of quantum gravity. These predictions are particularly significant, as they can be tested using current and future observational facilities, such as LIGO/Virgo, the Event Horizon Telescope, and X-ray observatories.
@@ Acknowledgments
We would like to express our sincere gratitude to the anonymous referee for their constructive suggestions and comments. B. C. L. is grateful to Excellence project PřF UHK 2211/2023-2024 for the financial support.